# Quantum Rationale-Aware Graph Contrastive Learning for Jet Discrimination

### v1

@@ III-B5 Classifier Head
III-C Benchmarked Models
We developed two classical models (GNN and EQGNN), three quantum models (QGNN, EQGNN, and QCL), five hybrid classical-quantum models (CQCL, 3 QCGCL variants, and QRGCL with $RX+H$ encoding) (see Table III), and classical counterpart of QRGCL, i.e., CRGCL (see Table II).
III-C1 Classical RGCL (CRGCL)
The classical RG (CRG) of RGCL is a GNN estimator designed to generate node representations from graph-structured data, which acts as the counterpart of QRGCL. It consists of three graph convolutional layers (GCN) that reduce feature dimensionality from the input to 32, 16, and finally, 8. Each layer is followed by batch normalization to stabilize training and reduce internal covariate shifts. ReLU activation is applied after the first two layers, while dropout is used to mitigate overfitting. The final output passes through a linear layer to yield a single value per node, followed by softmax for probabilistic representation. For a fair comparison between the CRG and QRG, we use the well-established ParticleNet as the encoder network, followed by the same projection head used in QRGCL.
III-C2 GNN
III-C3 Equivariant GNN (EGNN)
III-C4 Quantum GNN (QGNN)
III-C5 Equivariant Quantum GNN (EQGNN)
EQGNNs [15] are quantum analogs of EGNNs, incorporating equivariance into the quantum architecture. Like QGNNs, EQGNNs operate on quantum states but ensure that the learned representations are equivariant under symmetries, such as permutations or rotations. The final quantum states are aggregated through pooling, ensuring the network remains permutation-equivariant. This aggregation is followed by classical post-processing to yield the model’s predictions.
III-C6 QCL
QCL employs a quantum convolutional neural network (QCNN) with input size $18\times 18$ with zero padding as the encoder, where data-reuploading circuits (DRCs) replace classical convolutional kernels. The QCNN consists of 2 similar layers: the first with a ($3\times 3$) kernel, stride 2, and 3-qubit DRC as filter. The final encoding size after flattening is 16. The projection network either uses a 2-node linear layer.
III-C7 Classical-Quantum CL (CQCL)
III-C8 Quantum-Classical GCL (QCGCL)
QCGCL architecture initially uses multiple layers of graph attention-based convolution (GATConv) layers to capture graph features. Each GAT layer is followed by batch normalization and residual connections, which allow the model to skip layers for improved training. The graph-level embedding is achieved through mean and max pooling. A quantum circuit then processes this pooled GNN output. We benchmarked QCGCL using 3 different quantum encoders, as depicted in Table III. Finally, the GNN and quantum outputs are concatenated and passed through a fully connected readout layer to generate the final prediction.
TABLE I: Hyperparameter optimization results of the proposed QRGCL
Parameter type Parameter Test Accuracy AUC F1 Score
@@ III-C8 Quantum-Classical GCL (QCGCL)
TABLE II: Comparison between Classical and Quantum RG of RGCL.
@@
Test Accuracy AUC F1-score Test Accuracy AUC F1-score
@@
TABLE III: Performance benchmarking and ablation test of the proposed QRGCL. Bold indicates the best performance.
Model Test accuracy AUC F1 score # Parameters $n_{\alpha}$ $n_{layer}$ Batch size Epoch Encoder
@@
IV Experimental Setup
IV-A Simulation Tools and Environment
IV-B Hyperparameters and Configurations
The choice of epochs and batch sizes varies based on the computational requirements of classical and quantum models. Classical models like GNN and EGNN use larger batch sizes (64) and fewer epochs (19) to leverage efficient gradient updates, allowing faster convergence. In contrast, quantum models such as QGNN and EQGNN use much smaller batch sizes (1) due to quantum hardware limitations yet still require 19 epochs to ensure adequate learning despite more minor updates per step. For fully quantum and hybrid models like QRGCL and QCGCL, larger batch sizes (2000) and more epochs (50) are used due to the longer backpropagation caused by the complex custom loss function, with multiple loss components ensuring that the model has sufficient time to learn robust quantum-based representations, as the slower convergence can hinder effective learning.
V Results and Discussion
@@
The hyperparameters specific to QRGCL include encoder type, learning rate, and entanglement type, as detailed in Table I. Among the encoders, $RX$ and Hadamard ($H$) achieved the highest AUC values at 76.30% and 76.71%, respectively, while displacement-amplitude and $RZ$ encodings performed poorly, with AUCs of 72.10% and 67.32%. The optimal learning rate of $1\times 10^{-3}$ produced the highest AUC of 76.20%, with higher rates resulting in decreased performance. The $SWAP$ entanglement type yielded the best overall results, achieving an AUC of 76.20%. $CNOT$ and $CZ$ entanglements performed strongly, while other configurations underperformed. For the proposed QGRCL, two possible combinations of $RX$ and $H$ encoders the learning rate of $1\times 10^{-3}$ and $SWAP$ entangler were tried.
For both the classical RGCL and QRGCL models, tunable parameters included the number of nodes per head ($n_{\alpha}$) ranging from 7 to 10, the number of classical or quantum layers ($n_{layer}$) ranging from 2 to 5, and the augmentation ratio ranging from 0.1 to 0.3. As shown in Table II, QRGCL outperforms the classical RGCL with 8 and 10 nodes, achieving higher AUC values. At 3 layers, QRGCL surpasses the classical model, achieving 74.71% compared to 73.58%. However, it underperformed at 2 and 4 layers. With a 0.1 augmentation ratio, QRGCL significantly outperformed the classical approach, achieving 78.78% compared to 74.09%. Overall, the optimal parameters for QRGCL were found to be $n_{\alpha}=7$, $n_{layer}=3$, and an augmentation ratio of 10%.
Based on Table III and Figure 5b, it is evident that our proposed QRGCL model achieves the highest AUC score (77.53%). We took the best two well-performing encoders from Table I and developed two variants hybridizing them, and further experimentations revealed that $H$ followed by $RX$ gate is more robust. Additionally, the relatively large number of parameters in the QRGCL is due to the utilization of the ParticleNet GNN encoder, which possesses 125k parameters, while the QRG circuit only has 45 parameters. Tables I, II, and III complement the ablation studies for QRGCL by presenting results for configurations without rationale-awareness (QGNN, GNN, EGNN, EQGNN), analyzing VQC components (such as encoding variants, entanglement structures, and variations in qubits and layers), exploring different classical-quantum interfaces (including quantum-only and hybrid architectures, as well as a classical-only baseline), and examining the effects of rationale-guided data augmentation. Figure 5a shows both training and testing losses decrease steadily in QRGCL, indicating practical training with a continuous reduction in error. By around 800 epochs, the losses stabilize, suggesting that the model has reached a convergence point where further training yields minimal improvements. The small gap between the two loss curves suggests minimal overfitting and generalization of test data. Both training and testing accuracies rise sharply within the first 200 epochs, reflecting rapid learning and an effective initial adjustment of model parameters. Around 800 epochs, both accuracies stabilize. The test accuracy levels off slightly above 70%, while training accuracy remains close but slightly lower. This stability indicates that QRGCL has reached its maximum performance capacity on this dataset, with limited fluctuations.
VI Conclusion

### v2

@@ 3.2.5 Classifier Head
3.3 Benchmark Models
We developed two classical models (GNN and EQGNN), three quantum models (QGNN, EQGNN, and QCL), five hybrid classical-quantum models (CQCL, 3 QCGCL variants, and QRGCL with $RX+H$ encoding) (see Table 3), and the classical counterpart of QRGCL, i.e., CRGCL (see Table 2). More details are shown in Appendix D.
Table 1: Hyperparameter optimization results of the proposed QRGCL. Bold indicates the best performance.
Parameter type Parameter Test Accuracy ($\uparrow$) AUC ($\uparrow$) F1 Score ($\uparrow$)
@@ 3.3 Benchmark Models
Table 2: Comparison between classical and quantum RG of RGCL. Bold indicates the best performance.
@@
Test Accuracy ($\uparrow$) AUC ($\uparrow$) F1-score ($\uparrow$) Test Accuracy ($\uparrow$) AUC ($\uparrow$) F1-score ($\uparrow$)
@@
Table 3: Performance benchmarking and ablation test of the proposed QRGCL. Bold indicates the best performance.
Model Test accuracy ($\uparrow$) AUC ($\uparrow$) F1 score ($\uparrow$) # Parameters ($\downarrow$) $n_{\alpha}$ ($\downarrow$) $n_{layer}$ ($\downarrow$) Batch size Epoch Encoder
@@
4 Experimental Setup
In our experiments, we utilized various tools and frameworks to implement and train the models. The classical models were implemented using PyTorch, while the quantum components were handled using PennyLane and Qiskit. To simulate graph-based operations, we used the Deep Graph Library (DGL). The full experimental setup, including hardware specifications, hyperparameter configurations, and detailed experimental procedures, can be found in Appendix E.
5 Results and Discussion
@@
QRGCL achieved the highest AUC (77.53%), outperforming all baselines. The best performance was obtained using $H$ followed by $RX$ encoding, a learning rate of $1\times 10^{-3}$, and $SWAP$ entanglement, as detailed in Table 1. As shown in Table 2, QRGCL outperformed classical RGCL in most settings, particularly with 3 layers, 7-node rationales, and an augmentation ratio of 10%. More detailed results are outlined in Appendix F.
We took the best two well-performing encoders from Table 1 and developed two variants by hybridizing them, and further experimentations revealed that $H$ followed by $RX$ gate is more robust. Additionally, the relatively large number of parameters in the QRGCL is due to the utilization of the ParticleNet GNN encoder, which possesses 125k parameters, while the QRG circuit only has 45 parameters. Tables 1, 2, and 3 complement the ablation studies for QRGCL. Full ablation results including encoder comparisons, entanglement strategies, layer and rationale variations, and rationale-aware vs. rationale-agnostic models are detailed in Appendix F.
6 Broader Impacts
QRGCL enables efficient, low-supervision feature extraction that can accelerate particle physics discoveries, inspire quantum-augmented ML methods, raise ethical and interpretability considerations, and generalize to broader graph-structured problems across science and industry.
7 Conclusion